# A weak regularity lemma for polynomials

### v1

@@ Abstract
Applications include strong bounds for a problem of Karam on generalized rank, as well as a new method to obtain upper bounds for fan-in parameters in arithmetic circuits. For example, we show that if the image of a polynomial map $\mathbf{P}\colon\mathbb{F}^{n}\to\mathbb{F}^{m}$ of degree $d$ does not contain a line, then $\mathbf{P}$ can be computed by a depth-$4$ homogeneous formula with bottom fan-in at most $d/2$ and top fan-in at most $(2m)^{C(d)}$ (with $C(d)=2^{(1+o(1))d}$). Such a result was previously known only with at least a tower-type bound on the top fan-in. One implication of our work is a certain “barrier” to arithmetic circuit lower bounds, in terms of the smallest degree of a polynomial curve contained in the image of the given polynomial map.
@@ 1 Introduction
Regularity lemmas for polynomials have served as a key component in a variety of important results, as in the proof of the the finite-field Gowers Inverse conjecture in additive combinatorics and higher-order Fourier analysis (e.g. [10, 29]), Stillman’s conjecture in commutative algebra [1], and the Reed-Muller weight distribution problem in coding theory [2, 17], to name but a few. They serve as a key component in higher-order arithmetic regularity lemmas (e.g., [8, 9, 10, 12, 16]). [^1] Roughly speaking, arithmetic regularity lemmas approximate general functions in terms of low-degree polynomials that are approximately independent. (This approximate independence is indispensable; to quote Green [9]: second-order arithmetic regularity lemmas “reduce the study of general functions… to the study of projections… onto quadratic factors. This, however, is of little consequence unless we can study those supposedly simple objects.”) As is usually the case with regularity lemmas, they unfortunately only have weak bounds, and therefore so does any result that relies on them. In this paper we introduce a weaker regularity lemma for polynomials, with far better bounds, that is nevertheless powerful enough for some applications. For example, we obtain strong bounds for a result of Karam [15] on the relation between the generalized rank of a polynomial and its range. Our results also yield a method to prove upper bounds for arithmetic circuits.
@@
Roughly speaking, the usual regularity lemmas for polynomials (see, e.g., Proposition 8.1 in [6], Lemma 2.4 in [10], or Lemma 3.11 in [9]) can be described as results about polynomial decompositions. Namely, they express any $n$-variate polynomial $P\colon\mathbb{F}^{n}\to\mathbb{F}$—and more generally, any polynomial map $\mathbf{P}\colon\mathbb{F}^{n}\to\mathbb{F}^{m}$—in terms of a bounded number (independent of $n$) of polynomials $\mathbf{X}=(X_{1},...,X_{k})$, with the property that the latter “behave” approximately like $k$ (jointly independent) variables. Although the precise definition of “behave” can differ depending on the context, the intuition is as follows: for many purposes, any polynomial can be analyzed as though it had only $k$ variables. The number $k$ of parts $X_{1},\ldots,X_{k}$ in the decomposition depends only on the degree $d$ and the number $m$ of the given polynomials $\mathbf{P}=(P_{1},\ldots,P_{m})$, and as mentioned above, this dependency—befitting a regularity lemma—can be quite astronomical, often having worse than a tower-type growth (see, e.g., the Remark following Lemma 2.4 in [10]). [^2] Interestingly, recent regularization results of Lampert and Ziegler [23, 21] give strong bounds; however, unlike regularity lemmas, these do not provide a decomposition of $\mathbf{P}$ as a polynomial in $X_{1},\ldots,X_{k}$, but rather place $\mathbf{P}$ in the ideal generated by them. The main reason for these weak bounds is that in order to guarantee approximate independence, the polynomial parts $X_{1},\ldots,X_{k}$, as well as all their linear combinations, are required to have rank (see Section 2) that is large as a function of $k$, their number. Thus, the larger the number of parts, the more difficult it is to satisfy the rank requirement, which translates to an extremely large number of steps in the process to construct them. [^3] See, e.g., Proposition 8.1 in [6] for a clean example of this phenomenon, where the only thing that guarantees that the process at all terminates is a small decrease in the lexicographic order of the tuple of degrees of $\mathbf{X}$ in each step.
@@
In the setting of graphs, Frieze and Kannan [7] famously obtained a weak regularity lemma that has both strong bounds and multiple interesting applications. For comparison, the usual graph regularity lemma of Szemerédi decomposes any graph—by partitioning its vertices—into a bounded number of parts (independent of the number of vertices) with the guarantee that almost all graph parts are “approximately” random. The weak graph regularity lemma decomposes any graph in the same sense, but comes with only one, global guarantee. Our weak regularity lemma for polynomials is partly inspired by this approach.
@@ 1.1 Weak regularity for polynomials
Intuitively, the benefit of having a weak regular decomposition for a polynomial map $\mathbf{P}$ is that it reveals some information about the univariate polynomials “induced” by $\mathbf{P}$. In particular, it enables us to argue about the image of $\mathbf{P}$ (for which standard rank-based tools are unsuitable, as they typically give information only about equidistribution); indeed, a corollary of the weak regularity lemma will be that any line that $\mathbf{X}$ significantly intersects must in fact be fully contained in the image of $\mathbf{X}$. This translates to a polynomial curve (univariate polynomial map) contained in the image of $\mathbf{P}$. See below for more about applications.
We note that throughout the paper we work with finite fields $\mathbb{F}$, for which the notion of weak regularity can be most elementarily defined. Moreover, rank-based methods, which we rely on extensively in the proof (as do proofs of other regularity lemmas for polynomials), tend to be most challenging in the finite-field setting. To get an analogue of our weak regularity lemma for infinite fields would require replacing the probabilistic notion of approximate independence in the definition of weak regularity below, for example by a geometric or by a commutative algebra analogue (see, e.g., [22, 6]), and adapting the proof accordingly. [^4] Technically, this would require replacing the notion of bias of a polynomial (see Section 2.2), which is defined for finite fields, by standard analogues for infinite fields. For the structure-vs-randomness theorem we rely on (Theorem 2.10), a geometric analogue over infinite fields has long been known [19, 28] (in fact, with a slightly better, linear bound).
@@
A tuple of polynomials $\mathbf{P}$ over $\mathbb{F}=\mathbb{F}_{q}$ has a weak $\epsilon$-regular decomposition of size $k$ if $\mathbf{P}\subseteq\mathbb{F}[\mathbf{X}]$ for some $k$ homogeneous polynomials $\mathbf{X}=(X_{1},\ldots,X_{k})$, such that:
@@
Note that the two additional additional conditions at the end of the definition are necessary to avoid trivialities. This is clearly the case for the first, which guarantees that the decomposition of $\mathbf{P}$ depends on $X_{1}$. As for the other additional condition, that $X_{1}$ is of maximal degree—note that we can decompose any polynomial $P=P(x_{1},\ldots,x_{n})$, of degree $d$ in $x_{1}$ say, as $\sum_{i=0}^{d}x_{1}^{i}C_{i}(x_{2},\ldots,x_{n})$; that is, we always have the small decomposition $P=F(x_{1},C_{0},\ldots,C_{d})$ (where $F(z,y_{0},y_{1},\ldots,y_{d})=\sum_{i=0}^{d}z^{i}y_{i}$). Although $x_{1}$ is genuinely independent of the other parts of the decomposition $C_{0},\ldots,C_{d}$, its degree $\deg(x_{1})=1$ is generally far from the largest degree among them.
Let us also mention two extreme cases in the definition of weak regularity. First, if the polynomial map $\mathbf{P}\colon\mathbb{F}^{n}\to\mathbb{F}^{m}$ is itself equidistributed, that is, $|\Pr[\mathbf{X}=y]-q^{-1}|\leq\epsilon q^{-m}$ for every $y\in\mathbb{F}^{m}$, then $\mathbf{P}$ trivially has a weak $\epsilon$-regular decomposition of size $m$ (with $\mathbf{X}=\mathbf{P}$). Second, every polynomial map $\mathbf{P}\colon\mathbb{F}^{n}\to\mathbb{F}^{m}$ has the trivial decomposition $\mathbf{P}=\mathbf{P}(\mathbf{X})$ with $\mathbf{X}=(x_{1},\ldots,x_{n})$ its variables, which is a weak 0-regular decomposition (assuming without loss of generality that $\mathbf{P}$ depends on $x_{1}$), though of a trivially large size $n$.
@@ 1.2 Applications
Is there an algebraic explanation for non-surjectivity of (multivariate) polynomials? An important class of non-surjective polynomials are those of the form $P=F(Q)$ with $1<\deg(Q)<\deg(P)$—called uni-multivariate decomposable [11]—for some non-surjective univariate $F$ (for example, $P=(Q(x_{1},...,x_{n}))^{t}$ with $\gcd(t,|\mathbb{F}|-1)>1$). For such a polynomial of degree $d$, it must be that $\deg(Q)\leq d/2$ (as $\deg(F)\geq 2$). We show that this is always the case, in the sense that every non-surjective polynomial can be written in terms of a rather small number of polynomials all of degree at most $d/2$ (despite most polynomials not being uni-multivariate decomposable [30]). In fact, we show a significant generalization of this to polynomial maps. Note that the image of a polynomial map holds information about the dependencies between its various polynomials.
Following Green and Tao ([10], Definition 1.6), let $\operatorname{rank}_{t}(P)$, for a polynomial $P$ and $t\geq 0$, denote the least number of polynomials of degree at most $t$ such that $P$ can be expressed as a function of only them. Note that this extends naturally to polynomial maps $\mathbf{P}$. One easy-to-state application of the weak regularity lemma is that any polynomial that is not surjective, and more generally any polynomial map $\mathbf{P}\colon\mathbb{F}^{n}\to\mathbb{F}^{m}$ whose image $\operatorname{Im}\mathbf{P}$ ($\subseteq\mathbb{F}^{m}$) does not contain a line must have low generalized rank $\operatorname{rank}_{t}(\mathbf{P})$, for some $t$ bounded away from $\deg(\mathbf{P}):=\max_{i}\deg(P_{i})$.
@@
We note that $\operatorname{udeg}(P)\leq\deg(P)$, and that $\operatorname{udeg}(\mathbf{P})=1$ if and only if $\operatorname{Im}\mathbf{P}$ contains a line; see Section 3.1 for more properties of univariate degree. As an example, we have $\operatorname{udeg}(Q^{t})\leq t$, for any polynomial $Q$, with equality for any surjective $Q$ (in which case $\operatorname{Im}(Q^{t})=\{x^{t}|x\in\mathbb{F}\}$) and some choices of $t$ that depends on $|\mathbb{F}|$. Note that in this last example $P=Q^{t}$, we have $\deg(P)/\operatorname{udeg}(P)=\deg(Q)$, and if we take $Q$ to be a random polynomial $Q\colon\mathbb{F}^{n}\to\mathbb{F}$ then $\operatorname{rk}_{\deg(Q)-1}(P)$ is unbounded (as $n\to\infty$; see Example 1.4 in [15]). This shows that we cannot hope to bound $\operatorname{rk}_{t}(P)$ for $t<\deg(P)/\operatorname{udeg}(P)$.
Karam [15] proved that $\operatorname{rank}_{t}(P)$, for $t=\deg(P)/\operatorname{udeg}(P)$, is, in fact, bounded (independently of nn). [^5] Karam’s result actually bounds the closely related notion of $\operatorname{rk}_{t}$; see Section 3 below.
Theorem 1.1 (Karam, Theorem 1.3 in [15]).
@@
The proof in [15] uses a variant of the regularity lemma for polynomials, and thus yields an extremely weak bound $\gamma(p,d)$ on $\operatorname{rk}_{t}(P)$.
@@
Roughly speaking, the proof follows by showing that for any weak $\epsilon$-regular decomposition $\mathbf{P}=\mathbf{F}(\mathbf{X})$, for $\epsilon$ bounded away from $1$, we have $\operatorname{udeg}(\mathbf{P})\leq\deg_{1}(\mathbf{F})$ (that is, $\operatorname{Im}\mathbf{U}\subseteq\operatorname{Im}\mathbf{P}$ for some non-constant curve $\mathbf{U}$ with $\deg(\mathbf{U})=\deg_{1}(\mathbf{F})$). Using the guarantees of the homogeneity of $\mathbf{X}$ and of the maximality of $\deg(X_{1})$ ($\deg(\mathbf{X})$), we will deduce $\deg_{1}(\mathbf{F})\leq\deg(\mathbf{P})/\deg(\mathbf{X})$, that is, $\deg(\mathbf{X})\leq\deg(\mathbf{P})/\operatorname{udeg}(\mathbf{P})=t$, which would allow us to bound $\operatorname{rk}_{t}$.
Finally, this machinery also has a direct implication to the study of arithmetic formulas in computational complexity theory, namely, depth-$4$ arithmetic formulas. A $\sum^{[r]}\prod^{[a]}\sum\prod^{[b]}$-formula (e.g., [18, 20]) for a polynomial $P$ is a decomposition of the form $P=\sum_{i=1}^{r}\prod_{j=1}^{a_{i}}Q_{i,j}$ with $a_{i}\leq a$ and $\deg(Q_{i,j})\leq b$. More generally, for a tuple of polynomials $\mathbf{P}=(P_{1},\ldots,P_{m})$, such a formula means $P_{k}\in\operatorname{Sp}\{\prod_{j=1}^{a_{i}}Q_{i,j}|1\leq i\leq r\}$ for all $k$. It is an important challenge to understand the interplay between any of the different “fan-in” parameters $(r,a,b)$. Trivially, if the bottom fan-in $b$ is allowed to be as large as $\deg(P)$, then $P$ has a $\sum^{[r]}\prod^{[a]}\sum\prod^{[b]}$-formula with $r=a=1$. The question becomes interesting when the bottom fan-in $b$ is bounded away from $\deg(P)$. Using the weak regularity lemma, we deduce an upper bound on the top fan-in $r$ from a lower bound on the univariate degree. [^6] It is worth noting that the model of depth-$4$ homogeneous arithmetic formulas with top fan-in $r$ is already powerful for $r=\Theta(\log n)$; see [20].
@@
Let $P\colon\mathbb{F}^{n}\to\mathbb{F}$ with $\deg(P)\leq d$, over any finite field $\mathbb{F}$ with $d<\operatorname{char}(\mathbb{F})$. Suppose $P\in\mathbb{F}[\mathbf{P}]$ with $\mathbf{P}=(P_{1},\ldots,P_{m})$, $\deg(\mathbf{P})\leq d$, such that $\operatorname{udeg}(\mathbf{P})\geq u$. Then there is a (homogeneous) $\sum^{[r]}\prod^{[u]}\sum\prod^{[d/u]}$-formula for $P$ with top fan-in $r\leq(2m)^{2^{d(1+o(1))}}$.
Contrapositively, this means that any strong enough lower bound method for the top fan-in of depth-$4$ arithmetic formulas (of a given bottom fan-in, size, etc.) must somehow avoid applying to polynomial maps of high univariate degree, and to the polynomials they generate. It therefore seems that the notion of univariate degree could be an important parameter to consider when studying questions about arithmetic circuits and their top fan-in, and for polynomial identity testing of depth-$4$ formulas with bounded fan-in (a topic of considerable interest, see e.g., [5]).
@@
Univariate degree should not be confused with the concept of an elusive function of Raz [27], which is a polynomial map $\mathbf{P}$ satisfying $\operatorname{Im}\mathbf{P}\nsubseteq\operatorname{Im}\mathbf{\Gamma}$, for every polynomial map $\mathbf{\Gamma}$ with given parameters. Indeed, univariate degree is about containment in the other direction: a lower bound on $\operatorname{udeg}(\mathbf{P})$ means that $\operatorname{Im}\mathbf{U}\nsubseteq\operatorname{Im}\mathbf{P}$, for every polynomial curve $\mathbf{U}$ of a given degree.

### v2

@@ Abstract
Applications include strong bounds for a problem of Karam on generalized rank, as well as a new method to obtain upper bounds for fan-in parameters in arithmetic circuits. For example, we show that if the image of a polynomial map $\mathbf{P}\colon\mathbb{F}^{n}\to\mathbb{F}^{m}$ of degree $d$ does not contain a line, then $\mathbf{P}$ can be computed by a depth-$4$ arithmetic formula with bottom fan-in at most $d/2$ and top fan-in at most $(2m)^{C(d)}$ (with $C(d)=2^{(1+o(1))d}$). One implication of our work is a certain “barrier” to arithmetic circuit lower bounds, in terms of the smallest degree of a polynomial curve contained in the image of the given polynomial map.
@@ 1 Introduction
Regularity lemmas for polynomials have served as a key component in a variety of important results, as in the proof of the the finite-field Gowers Inverse conjecture in additive combinatorics and higher-order Fourier analysis (e.g. [GT09, TZ12]), Stillman’s conjecture in commutative algebra [AH20], and the Reed-Muller weight distribution problem in coding theory [BL15, KLP12], to name but a few. They serve as a key component in higher-order arithmetic regularity lemmas (e.g., [GW11, Gree06, GT09, HL11, KL08]). [^1] Roughly speaking, arithmetic regularity lemmas approximate general functions in terms of low-degree polynomials that are approximately independent. (This approximate independence is indispensable; to quote Green [Gree06]: second-order arithmetic regularity lemmas “reduce the study of general functions… to the study of projections… onto quadratic factors. This, however, is of little consequence unless we can study those supposedly simple objects.”) As is usually the case with regularity lemmas, they unfortunately only have weak bounds, and therefore so does any result that relies on them. In this paper we introduce a weaker regularity lemma for polynomials, with far better bounds, that is nevertheless powerful enough for some applications. For example, we obtain strong bounds for a result of Karam [Kar23] on the relation between the generalized rank of a polynomial and its range. Our results also yield a method to prove upper bounds for arithmetic circuits.
@@
Roughly speaking, the usual regularity lemmas for polynomials (see, e.g., Proposition 8.1 in [ESS19], Lemma 2.4 in [GT09], or Lemma 3.11 in [Gree06]) can be described as results about polynomial decompositions. Namely, they express any $n$-variate polynomial $P\colon\mathbb{F}^{n}\to\mathbb{F}$—and more generally, any polynomial map $\mathbf{P}\colon\mathbb{F}^{n}\to\mathbb{F}^{m}$—as in terms of a bounded number (independent of $n$) of polynomials $\mathbf{X}=(X_{1},...,X_{k})$, that is, $P=F(\mathbf{X})$, with the property that the polynomials in $\mathbf{X}$ “behave” approximately like $k$ (jointly independent) variables. Although the precise definition of “behave” can differ depending on the context, the intuition is as follows: for many purposes, any polynomial can be analyzed as though it had only $k$ variables. The number $k$ of parts $X_{1},\ldots,X_{k}$ in the decomposition depends only on the degree $d$ and the number $m$ of the given polynomials $\mathbf{P}=(P_{1},\ldots,P_{m})$, and as mentioned above, this dependency—befitting a regularity lemma—can be quite astronomical, often having worse than a tower-type growth (see, e.g., the Remark following Lemma 2.4 in [GT09]). [^2] Interestingly, recent regularization results of Lampert and Ziegler [LZ24, Lam23] give strong bounds; however, unlike regularity lemmas, these do not provide a decomposition of $\mathbf{P}$ as a polynomial in $X_{1},\ldots,X_{k}$, but rather place $\mathbf{P}$ in the ideal generated by them. The main reason for these weak bounds is that in order to guarantee approximate independence, the polynomial parts $X_{1},\ldots,X_{k}$, as well as all their linear combinations, are required to have rank (see Section 2) that is large as a function of $k$, their number. Thus, the larger the number of parts, the more difficult it is to satisfy the rank requirement, which translates to an extremely large number of steps in the process to construct them. [^3] See, e.g., Proposition 8.1 in [ESS19] for a clean example of this phenomenon, where the only thing that guarantees that the process at all terminates is a small decrease in the lexicographic order of the tuple of degrees of $\mathbf{X}$ in each step.
@@
In the setting of graphs, Frieze and Kannan [FK99] famously obtained a weak regularity lemma that has both strong bounds and multiple interesting applications. For comparison, the usual graph regularity lemma of Szemerédi decomposes any graph—by partitioning its vertices—into a bounded number of parts (independent of the number of vertices) with the guarantee that almost all graph parts are “approximately” random. The weak graph regularity lemma decomposes any graph in the same sense, but comes with only one, global guarantee. Our weak regularity lemma for polynomials is partly inspired by this approach.
@@ 1.1 Weak regularity for polynomials
Intuitively, the benefit of having a weak regular decomposition for a polynomial map $\mathbf{P}$ is that it reveals some information about the univariate polynomials “induced” by $\mathbf{P}$. In particular, it enables us to argue about the image of $\mathbf{P}$ (for which standard rank-based tools are unsuitable, as they typically give information only about equidistribution); indeed, a corollary of the weak regularity lemma will be that any line that $\mathbf{X}$ intersects somewhat often must in fact be fully contained in the image of $\mathbf{X}$. This translates to a polynomial curve (univariate polynomial map) contained in the image of $\mathbf{P}$. See below for more about applications.
We note that throughout the paper we work with finite fields $\mathbb{F}$, for which the notion of weak regularity can be most elementarily defined. Moreover, rank-based methods, which we rely on extensively in the proof (as do proofs of other regularity lemmas for polynomials), tend to be most challenging in the finite-field setting. [^4] To get an analogue of our weak regularity lemma for infinite fields would require replacing the probabilistic notion of approximate independence in the definition of weak regularity below, for example by a geometric or by a commutative algebra analogue (see, e.g., [LZ24, ESS19]), and adapting the proof accordingly. Technically, this would require replacing the notion of bias of a polynomial (see Section 2.2) by standard analogues for infinite fields. For the structure-vs-randomness theorem we rely on (Theorem 2.10), a geometric analogue over infinite fields has long been known [KLP24, Sch85] with, in fact, a slightly better, linear bound.
@@
A tuple of polynomials $\mathbf{P}$ over $\mathbb{F}=\mathbb{F}_{q}$ has a weak $\epsilon$-regular decomposition of size $k$ if $\mathbf{P}\subseteq\mathbb{F}[\mathbf{X}]$, for some $k$ homogeneous polynomials $\mathbf{X}=(X_{1},\ldots,X_{k})$, such that:
@@
Note that the two additional conditions at the end of the definition are necessary to avoid trivialities. This is clearly the case for the first, which guarantees that the decomposition of $\mathbf{P}$ depends on $X_{1}$. As for the other additional condition, that $X_{1}$ is of maximal degree—note that we can decompose any polynomial $P=P(x_{1},\ldots,x_{n})$, of degree $d$ in $x_{1}$ say, as $\sum_{i=0}^{d}x_{1}^{i}C_{i}(x_{2},\ldots,x_{n})$; that is, we always have a small decomposition $P=F(x_{1},C_{0},\ldots,C_{d})$ (where $F(z,y_{0},y_{1},\ldots,y_{d})=\sum_{i=0}^{d}z^{i}y_{i}$). Although $x_{1}$ is genuinely independent of the other parts of the decomposition $C_{0},\ldots,C_{d}$, its degree $\deg(x_{1})=1$ is generally far from the largest degree among them.
Let us also mention two extreme cases in the definition of weak regularity. First, if the polynomial tuple $\mathbf{P}$, viewed as a map $\mathbf{P}\colon\mathbb{F}^{n}\to\mathbb{F}^{m}$, is itself equidistributed, that is, $|\Pr[\mathbf{P}=y]-q^{-1}|\leq\epsilon q^{-m}$ for every $y\in\mathbb{F}^{m}$, then $\mathbf{P}$ trivially has a small weak $\epsilon$-regular decomposition (with $\mathbf{X}=\mathbf{P}$). Second, $\mathbf{P}$ always has the trivial decomposition $\mathbf{P}=\mathbf{P}(\mathbf{X})$ with $\mathbf{X}=(x_{1},\ldots,x_{n})$ its variables, which is a weak 0-regular (i.e., a perfectly weak regular decomposition; assuming without loss of generality that $\mathbf{P}$ depends on $x_{1}$), though of a trivially large size $n$.
@@ 1.2 Applications
Is there an algebraic explanation for non-surjectivity of (multivariate) polynomials? An important class of non-surjective polynomials are those of the form $P=F(Q)$ with $1<\deg(Q)<\deg(P)$—sometimes called uni-multivariate decomposable [GRS02]—for some non-surjective univariate $F$ (for example, $P=(Q(x_{1},...,x_{n}))^{t}$ with $\gcd(t,|\mathbb{F}|-1)>1$). For such a polynomial of degree $d$, it must be that $\deg(Q)\leq d/2$ (as $\deg(F)\geq 2$). We show that this is always the case, in the sense that every non-surjective polynomial can be written in terms of a rather small number of polynomials all of degree at most $d/2$ (despite most polynomials not being uni-multivariate decomposable [VZ15]). In fact, we show a significant generalization of this to polynomial maps. Note that the image of a polynomial map holds information about the dependencies between its various polynomials.
Following Green and Tao ([GT09], Definition 1.6), let $\operatorname{rank}_{t}(P)$, for a polynomial $P$ and $t\geq 0$, denote the least number of polynomials of degree at most $t$ such that $P$ can be expressed as a function of only them. Note that this extends naturally to polynomial maps $\mathbf{P}$. One easy-to-state application of the weak regularity lemma is that any polynomial that is not surjective, and more generally any polynomial map $\mathbf{P}\colon\mathbb{F}^{n}\to\mathbb{F}^{m}$ whose image $\operatorname{Im}\mathbf{P}$ ($\subseteq\mathbb{F}^{m}$) does not contain a line, must have low generalized rank $\operatorname{rank}_{t}(\mathbf{P})$, for some $t$ bounded away from $\deg(\mathbf{P}):=\max_{i}\deg(P_{i})$.
@@
We note that $\operatorname{udeg}(P)\leq\deg(P)$, and that $\operatorname{udeg}(\mathbf{P})=1$ if and only if $\operatorname{Im}\mathbf{P}$ contains a line; see Section 3.1 for more properties of univariate degree. As an example, we have $\operatorname{udeg}(Q^{t})\leq t$, for any polynomial $Q$, with equality for any surjective $Q$ (in which case $\operatorname{Im}(Q^{t})=\{x^{t}|x\in\mathbb{F}\}$) and some choices of $t$ that depends on $|\mathbb{F}|$. Note that in this last example $P=Q^{t}$, we have $\deg(P)/\operatorname{udeg}(P)=\deg(Q)$, and if we take $Q$ to be a random polynomial $Q\colon\mathbb{F}^{n}\to\mathbb{F}$ then $\operatorname{rk}_{\deg(Q)-1}(P)$ is unbounded (as $n\to\infty$; see Example 1.4 in [Kar23]). This shows that we cannot hope to bound $\operatorname{rk}_{t}(P)$ for $t<\deg(P)/\operatorname{udeg}(P)$.
Karam [Kar23] proved that $\operatorname{rank}_{t}(P)$, for $t=\deg(P)/\operatorname{udeg}(P)$, is, in fact, bounded (independently of nn). [^5] Karam’s result actually bounds the closely related notion of $\operatorname{rk}_{t}$; see Section 3 below.
Theorem 1.1 (Karam, Theorem 1.3 in [Kar23]).
@@
The proof in [Kar23] uses a variant of the regularity lemma for polynomials, and thus yields an extremely weak bound $\gamma(p,d)$ on $\operatorname{rk}_{t}(P)$.
@@
Roughly speaking, the proof follows by showing that for any weak $\epsilon$-regular decomposition $\mathbf{P}=\mathbf{F}(\mathbf{X})$, for $\epsilon$ bounded away from $1$, we have $\operatorname{udeg}(\mathbf{P})\leq\deg_{1}(\mathbf{F})$, the degree of $\mathbf{F}$ in its first variable (so $\operatorname{Im}\mathbf{U}\subseteq\operatorname{Im}\mathbf{P}$ for some non-constant curve $\mathbf{U}$ with $\deg(\mathbf{U})=\deg_{1}(\mathbf{F})$). Using the guarantees of the homogeneity of $\mathbf{X}$ and of the maximality of $\deg(X_{1})$ ($\deg(\mathbf{X})$), we will deduce $\deg_{1}(\mathbf{F})\leq\deg(\mathbf{P})/\deg(\mathbf{X})$, that is, $\deg(\mathbf{X})\leq\deg(\mathbf{P})/\operatorname{udeg}(\mathbf{P})=t$, which would allow us to bound $\operatorname{rk}_{t}$.
Finally, this machinery also has a direct implication to the study of arithmetic formulas in computational complexity theory, namely, depth-$4$ arithmetic formulas. A $\sum^{[r]}\prod^{[a]}\sum\prod^{[b]}$-formula (e.g., [KSS14, KS14]) for a polynomial $P$ is a decomposition of the form $P=\sum_{i=1}^{r}\prod_{j=1}^{a_{i}}Q_{i,j}$ with $a_{i}\leq a$ and $\deg(Q_{i,j})\leq b$. It is an important challenge to understand the interplay between any of the different “fan-in” parameters $(r,a,b)$. Trivially, if the bottom fan-in $b$ is allowed to be as large as $\deg(P)$, then $P$ has a $\sum^{[r]}\prod^{[a]}\sum\prod^{[b]}$-formula with $r=a=1$. The question becomes interesting when the bottom fan-in $b$ is bounded away from $\deg(P)$. Using the weak regularity lemma, we deduce an upper bound on the top fan-in $r$ from a lower bound on the univariate degree. [^6] It is worth noting that the model of depth-$4$ homogeneous arithmetic formulas with top fan-in $r$ is already powerful for $r=\Theta(\log n)$; see [KS14].
@@
Let $P\colon\mathbb{F}^{n}\to\mathbb{F}$ with $\deg(P)\leq d$, over any finite field $\mathbb{F}$ with $d<\operatorname{char}(\mathbb{F})$. Suppose $P\in\mathbb{F}[\mathbf{P}]$ with $\mathbf{P}=(P_{1},\ldots,P_{m})$, $\deg(\mathbf{P})\leq d$, such that $\operatorname{udeg}(\mathbf{P})\geq u$. Then there is a $\sum^{[r]}\prod^{[2u]}\sum\prod^{[d/u]}$-formula for $P$ with top fan-in $r\leq(2m)^{2^{d(1+o(1))}}$. Moreover, the formula is homogeneous [^7] That is, each $Q_{i,j}$ is homogeneous, and all products $\prod_{j}Q_{i,j}$ are of the same degree. if $P$ is.
Contrapositively, this means that any strong enough lower bound method for the top fan-in of depth-$4$ arithmetic formulas (of a given bottom fan-in, size, etc.) must somehow avoid applying to polynomial maps of high univariate degree, and to the polynomials they generate. It therefore seems that the notion of univariate degree could be an important parameter to consider when studying questions about arithmetic circuits and their top fan-in, and for polynomial identity testing of depth-$4$ formulas with bounded fan-in (a topic of considerable interest, see e.g., [DDS21]).
@@
Univariate degree should not be confused with the concept of an elusive function of Raz [Raz10], which is a polynomial map $\mathbf{P}$ satisfying $\operatorname{Im}\mathbf{P}\nsubseteq\operatorname{Im}\mathbf{\Gamma}$, for every polynomial map $\mathbf{\Gamma}$ with given parameters. Indeed, univariate degree is about containment in the other direction: a lower bound on $\operatorname{udeg}(\mathbf{P})$ means that $\operatorname{Im}\mathbf{U}\nsubseteq\operatorname{Im}\mathbf{P}$, for every polynomial curve $\mathbf{U}$ of a given degree.